OpenStudy (anonymous):

What is 16% of 40? A. 3.2 B. 16 C. 6.4 D. 8

OpenStudy (aroub):

C. 6.4

OpenStudy (anonymous):

C

OpenStudy (anonymous):

C

myininaya (myininaya):

\[\frac{16}{100} \cdot 40=\frac{16}{10} \cdot 4 =\frac{8}{5} \cdot 4=\frac{32}{5}\]